OpenStudy (ac3):

Find the volume of the region bounded by the paraboloid z=X^(2)+y^(2) and below the triangle enclosed by the lines y=x,x=0,and x+y=2 in the xy-plane.

OpenStudy (ac3):

@greatlife44 have time to help me with this?

OpenStudy (ac3):

how much calculus do you know @Johan14th ?

OpenStudy (johan14th):

Sorry not really sure on this one :(

OpenStudy (ac3):

need to know multi-variable calculus to do this. That's where you learn how to do this problem. no worries

terenzreignz (terenzreignz):

This takes me back ^^ \[\Large \iint_R x^2+y^2dA\]

terenzreignz (terenzreignz):

@Ac3 Still here, are you? ^^

terenzreignz (terenzreignz):

Well, while you're still somewhere else, I might as well sketch a graph of the region R in question...

terenzreignz (terenzreignz):



OpenStudy (ac3):

ok good i set up the integral as....

OpenStudy (ac3):

\[\int\limits_{0}^{1}\int\limits_{x}^{2-x}x ^{2}+y ^{2}dydx\]

OpenStudy (ac3):

and when i solved i got (23/18) which is wrong answer is (4/3)

terenzreignz (terenzreignz):

Hang on, let's see: \[\Large = \int_0^1\left[x^2y + \frac13y^3\right]_{y=x}^{y=2-x}\]

OpenStudy (ac3):

(2-x)^3 expanded = 8-4x-8x+4x^(2)+2x^(2)-x^(3) you can double check me though

OpenStudy (ac3):

unless i set up the integral incorrectly

terenzreignz (terenzreignz):

dx. Oops. Anyway, this becomes \[\large = \int_0^1\left[x^2(2-x) + \frac13(2-x)^3\right]-\left[x^3 + \frac13x^3\right]dx\]

OpenStudy (ac3):

ok nice that's what i had

terenzreignz (terenzreignz):

I'm sorry ^^ I'm not usually this slow, but I haven't done this in a long while :D

OpenStudy (ac3):

no worries man so far so good you have exactly what i got only i got it wrong lol

OpenStudy (ac3):

i'm not sure if i set up that double integral properly

terenzreignz (terenzreignz):

Means I probably will too :( Oh well \[\large = \int_0^1\left[2x^2-x^3 + \frac13(8 -12x+6x^2-x^3)\right]-\left[x^3 + \frac13x^3\right]dx\]

terenzreignz (terenzreignz):

\[\large = \int_0^1\left[2x^2-x^3 + \frac83 -4x+2x^2-\frac{x^3}3\right]-\left[x^3 + \frac{x^3}3\right]dx\]

OpenStudy (ac3):

hmm keep going maybe you'll get the right answer

terenzreignz (terenzreignz):

Judging by that reaction, have I deviated from your solution?

OpenStudy (ac3):

so far yes

OpenStudy (ac3):

maybe i calculated it incorrectly

terenzreignz (terenzreignz):

In that case \[\Large \int_0^1 \left[-\frac83 x^3 + 4 x^2-4x+ \frac83\right]dx\]

OpenStudy (ac3):

YOU GOT IT!!!

OpenStudy (ac3):

ah dam it so i did set it up right i just have to make sure i simplify properly

terenzreignz (terenzreignz):

Should be \[\Large \left[-\frac23x^4 + \frac43x^3-2x^2 + \frac83x\right]_0^1\]

OpenStudy (ac3):

well done

terenzreignz (terenzreignz):

Looks like I'm not that rusty after all ^^ It has been years, you know XD

OpenStudy (ac3):

how do i give you a medal

terenzreignz (terenzreignz):

you just did ^^ Well... this was fun :D

terenzreignz (terenzreignz):

Don't you just HATE these problems which are so susceptible to human error? :D

OpenStudy (ac3):

oh so that's how you do it. Yes i hate them now i'm gonna re-do and make sure i get the right answer

OpenStudy (ac3):

it's like they hand-craft them on purpose

terenzreignz (terenzreignz):

Well, see ya around ^^